OpenStudy (anonymous):

Find the surface area of a cone with radius of 8.8 cm and slant height of 7.1 cm to the nearest square unit. Use 3.14 for pi.

OpenStudy (anonymous):

SA = πr^2 + πrl where l = slant height. Just plug the numbers into this formula.

OpenStudy (anonymous):

can you explain less mathematically\ how you would explain to a beigginer

OpenStudy (anonymous):

Ok. the formula I posted above is the formula for surface area for a cone. Just plug in numbers for the values they represent. SA = πr^2 + πrl SA = π(8.8)^2 + π(8.8)(7.1) Just multiply and add.

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

You're welcome :)